What is the mass of 2.234 x 10²³ (2.234E23) atoms of sulfur? (S = 32.07 g)
Round to the nearest tenth (1 decimal places)
1
Expert's answer
2020-03-23T10:14:38-0400

The amount in moles of the sulfur sample could be determined via the following:


"n(S)=\frac{N_{atoms}}{N_{A}};"


On the other hand, the amount in moles could be determined via the following:


"n(S)=\frac{m_{S}}{M_{S}};"


After combining of the both expressions:


"\frac{m_{S}}{M_{S}}=\frac{N_{atoms}}{N_{A}},"


or


"m_{S}=\frac{N*M_{S}}{N_{S}}=\frac{2.234*10^{23}*32.07g/mol}{6.022*10^{23}mol^{-1}}=11.90g \sim 11.9g."